# HAAP: Vision-context Hierarchical Attention Autoregressive with Adaptive Permutation for Scene Text Recognition

### v1

@@ IV-D Ablation study
TABLE VI: The ablation study of CHA.
PLM IPN MHA CHA number of IR (Avg-9.(%))
0 1 2
- - 89.81 89.96 90.12
- - 90.15 90.17 90.15
- - 89.97 90.21 90.28
- - 90.40 90.43 90.42
The qualitative results of IPN and CHA are shown in Fig. 9. We find that the introduction of IPN and CHA makes the model not only effective for arbitrary orientation character recognition but also has strong generalization ability for low-quality and occlusion.
@@ IV-E Limitation and discussion
Although the proposed method can accurately recognize complex texts of arbitrary shapes in most cases, it still struggles with four kinds of tiny texts (image pixels not exceeding 2000): i:) front and back view severely confused; ii:) front and back text sticking together caused by blurring over long distances; iii:) shape-shifting with multiple occlusions; and iv:) alternating overlapping. Fig. 10 shows examples of such failure cases. The model loses attention on details during learning as the representation of character features declines in sequences.
@@ V Conclusion
To improve the autoregressive generalization based on internal LM, this paper proposes HAAP to enhance the location-context-image interaction capability by designing IPN and CHA. First, IPN is used to parse sequence meaning and syntactic structure by dynamically specifying token dependencies with an adaptive attention mask. The adaptive AR subset increases the diversity of the training data and prevents model dependency on specific sequences. The training overhead of PLM can be reduced and training fitting oscillations is avoided. Second, CHA module for hierarchical feature processing is proposed to exploit position-context-image semantic dependencies in the sequence without IR. The experimental results validate the effectiveness and the SOTA performance of of HAAP in terms of accuracy, complexity, and latency. In the future, we expect to extend HAAP to end-to-end detection and recognition, where the detector can be embedded in the visual decoder since the strong region perception is exhibited using the self-attention mechanism.
@@ Acknowledgment
This work was supported by the Natural Science Fund of China under Grants 62171135, Industry-University Research Project of Education Department 2022.

### v2

@@ IV-D Ablation study
The effectiveness of CHA To evaluate the effectiveness of CHA, we compare PLM-based and IPN-based models with CHA in different numbers of IRs. As shown in Table. VII, we find that both the PLM model and IPN model without CHA require additional IR to improve the autoregressive generalization. By using CHA, the PLM model achieves the Avg9Avg-9 of $90.15\%$ without IR, outperforming the one after iterating IR three times without CHA ($90.14\%$). Based on CHA, the IPN model achieves a performance gain of $0.43\%$ without IR, indicating the effectiveness of CHA in decoding position tokens. In addition, Fig. 9 shows the trade-off between the Avg9Avg-9, the model computational complexity (FLOPs), and single-image latency (Latency) with the number of IRs. The model without CHA requires two times of IR to achieve the stable performance, while the one with CHA maintains the performance advantage regardless of IR. Thus, the use of CHA avoids the additional computational complexity and latency induced by IR.
As shown in Fig. 10, although IR enhances the confidence of the IPN model, it results in wrong predictions. In contrast, the IPN model with CHA can generate high-confidence and accurate predictions without IR. The results clearly show that CHA can improve the performance of the autoregressive model through position-context-image interactions without the need for additional contextual semantic aggregation through IR. The qualitative results in Fig. 11 show that the CHA-based model effectively recognizes low-resolution and blurred texts.
TABLE VII: The ablation study of CHA.
Algorithm and complexity Number of IR (Avg-9.(%))
0 1 2 3 4
PLM w/o CHA 89.81 89.96 90.12 90.14 90.12
IPN w/o CHA 89.97 90.21 90.28 90.30 90.31
FLOPs(G) 3.38 3.54 3.70 3.86 4.02
PLM with CHA 90.15 90.17 90.15 90.16 90.13
IPN with CHA 90.40 90.43 90.42 90.42 90.41
FLOPs(G) 3.62 3.81 4.00 4.19 4.38
Figure 9: Average accuracy (Avg-9) in nine datasets and FLOPs of the model (Left) and single-image latency (Right) for each decoding scheme. The number of IR used is indicated for each point.
Figure 10: Comparison of confidence heat maps. Red color indicates the highest score, and blue color indicates the lowest score. For each approach from top to bottom are the score region visualization, confidence level, and prediction results. Ground truths are shown below the original images.
@@ IV-E Limitation and discussion
Although the proposed method can accurately recognize arbitrary-shaped texts in most cases, it still struggles with four kinds of tiny texts (the total number of pixels does not exceed 2000) due to the lack of visual representation of tiny text.: i:) target confused with the background; ii:) blurred long-text; iii:) deformed text under multi-occlusion; and iv:) overlapping text. Fig. 12 shows examples of such failure cases.
@@ V Conclusion
This paper proposes HAAP to enhance the position-context-image interaction capability using IPN and CHA to improve autoregressive generalization. First, IPN captures the correlation between context and visual information by dynamically specifying token dependencies with an adaptive attention mask. Adaptive sequence modeling increases the diversity of the training data and prevents the model from training fit oscillation. Second, CHA module for hierarchical feature processing is proposed to exploit position-context-image semantic dependencies in the sequence without IR. The experimental results validate the effectiveness and SOTA performance of HAAP in terms of accuracy, complexity, and latency. In the future, we expect to extend HAAP to end-to-end detection and recognition, where the detector can be embedded in the visual decoder since position awareness is exhibited using the attention mechanism.
@@ Acknowledgment
This work was supported by the NSFC (No. 62171135), Fujian Key Project (No.2023XQ004), Fujian Eagle-Scholar and Distinguished Project (No. 2022J06010), Fujian Natural Science Fund (No. 2024H6013).